Fund the volume generated by revolving the area enclosed by the curves x^2+y^2=36 about the y-axis.
1
Expert's answer
2020-12-21T18:11:11-0500

The curves x^2+y^2=36 make a circle with radius of 6. The figure generated by revolving the area is a sphere, so the volume of this sphere is:

"V = 4/3 πr³ = 4/3×\pi×6^3 = 904.78."